A random variable X has a normal distribution with mean 45 and variance of 9.



Find the range of scores that lie:




a. within one standard deviation from the mean



b. within two standard deviations from the mean




1
Expert's answer
2022-04-06T09:32:47-0400

"\text{The mean }\mu=45,\text{the standard deviation }\sigma=\sqrt{9}=3."


a. The range of scores that lie within one standard deviation from the mean :

"(\mu-\sigma,\mu +\sigma)=(45-3,45+3)=(42,48)."


b. The range of scores that lie within two standard deviations from the mean :

"(\mu-2\sigma,\mu +2\sigma)=(45-2\cdot3,45+2\cdot3)=(39,51)."